OpenStudy (anonymous):

The table below shows the number of hours some business people in two states spend in meetings each week: State A 21 23 24 22 24 25 23 23 22 State B 24 22 20 23 23 50 20 46 21 Part A: Create a five-number summary and calculate the interquartile range for the two sets of data. (6 points) Part B: Are the box plots symmetric? Justify your answer. (4 points)

OpenStudy (anonymous):

It will not let me view any of that..

OpenStudy (ybarrap):

Control-Click the links above

OpenStudy (anonymous):

I

OpenStudy (anonymous):

It still isn't showing

OpenStudy (perl):

*

OpenStudy (anonymous):

I still dont understand

OpenStudy (ybarrap):

Maybe this will help clarify. The interquartile range (IQR) is a measure of variability, based on dividing a data set into quartiles. Quartiles divide a rank-ordered data set into four equal parts. The values that divide each part are called the first, second, and third quartiles; and they are denoted by Q1, Q2, and Q3, respectively. Q1 is the "middle" value in the first half of the rank-ordered data set. Q2 is the median value in the set. Q3 is the "middle" value in the second half of the rank-ordered data set. The interquartile range is equal to Q3 minus Q1. Your data for State A in rank order: 21 22 22 Q1 23 23 Q2 23 24 Q3 24 25 So IQR is 24-22=2 5-number Summary Min = 20 Q1 = 21 Q2 = 23 Q3 = 24 Max = 50 Box Plot IS Symmetric:  For State B: 20 20 21 Q1 22 23 Q2 23 24 Q3 46 50 IQR = Q3-Q1=3 5-number Summary Min = 20 Q1 = 21 Q2 = 23 Q3 = 24 Max =50 Box Plot is NOT symmetric:  Does this make sense?